From mar@roe.ac.uk Wed Mar 5 09:25:14 2014 Date: Tue, 4 Mar 2014 18:13:30 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Dave Morris , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 28 February 2014 WFAU VDFS Science Archives progress meeting minutes, 28 February 2014 Present: RPB, MAR, ETWS, NJC, ACD Apologies: MSH, RGM, RSC, NCH, KTN, AL DoNM: 10am Friday 14 March 2014 in the Crawford Meeting Room MR2 ACD in the chair (room is booked). ** Note: CHANGE OF VENUE to MR2 Actions discharged this time: ---------------------------- ACTION: RPB to do the work necessary to make the pending VISTA public releases. Done. ACTION: RGM: E-mail Mike Irwin to ask for an example of the satellite data. Done, we had been given access to some example files. ACTION: NJC: Set up a schema to hold these data. Done, the files had been ingested and a database release made. ACTION: RPB to ask NJC to carry out checks on the latest VVV release when it was ready, in preparation for a release. Done, VVVDR2 has been released to the consortium. Actions partly discharged but continuing ---------------------------------------- ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Testing ongoing and Twiki page available for any additional suggestions Actions carried forward from 7/2/2014 meeting: ---------------------------------------------- ACTION: RSC to investigate modifying CU7 to recover ATLAS VST unmatched sources around frameset boundaries. Continues. Investigations had begun and a strategy had been worked out. However, this action now has lower priority. Nonetheless a fix should be put in place before the ATLAS meeting on 14-16 April and it should be followed by a consortium release shortly after the meeting. ACTION: NJC to perform a quick check on the VMC public release. Continues. ACTION: ALL to add/maintain entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Continues. ACTION: ACD to discuss with RGM the available budget for hardware purchases. Continues. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy (http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan). However, it needs more work and in particular RGM still had to determine which WFAU archives had cross-matches with which SDSS releases. ACTION: NJC to sort out the magnitude limits for V1.3 data. Continues. The software was largely written, but NJC had not had the time to finish and test. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: ALL: make sure that all forthcoming absences during the next couple of months are entered into the 'staff absences' Wiki page. Continues: needed for planning next WFAU quarterly ACTION: RSC to set a date for completing the actions from the brainstorming meeting held shortly before Christmas. (This action is in preparation for the next brainstorming meeting, which will be arranged once the VVV release is complete.) Continues. ACTION: NJC to remind Omar Almaini about the UKIDSS DR11 final data release. Continues. Specific points and new actions: ------------------------------- Project management: The new owners of UKIRT had visited on 21/2/14. The meeting went well with an example release database, based on the newly developed schema and the data provided by CASU, available for demonstration. Alex Ouzounis would be leaving WFAU for pastures new sometime in May. He will be missed, especially given the recent discovery that he can bake a really good cake! WFCAM, VISTA and VST updates: There's a possibility that we'll be the new home for the WFCAM raw data tape archive currently located in Hawaii. Comments and issues arising from CASU minutes: No new minutes this time. Networking: ETWS noted -- There will be a short UKLight downtime due to maintenance on 11/03/14. -- Transfers for VISTA Nov and Dec 2013 data have started. WSA/VSA/OSA/GES Operations: ETWS noted -- Created a new set of browser pages and metadocs for the latest VISTA public releases. -- VVV ingest are slowly nearing an end. Some broken files have been sent to CASU to be fixed. RPB noted that the VHS P90 release had finished, however room needed to freed up on the public servers in order to release it. Hardware and Systems: MSH reported that a new NAS box and curation/web server had been ordered. A test system, employing a new RAID card and set-up, would also be arriving. Survey Data Release: As mentioned earlier the latest VHS release would be released when space had been freed up on the public servers. VMC P91 release - QC had been applied and quality bit flagging and tile/pawprints could now be run. Software: ETWS noted -- Included new code in CU3 to ingest the new WFCAM satellite data. Similarly MAR reported making UI tweaks to accommodate the WFCAM satellite data release. NJC reported working on the code for comparing the zero-points and photometry of deeps and their constituent frames. He'd also worked on the WFCAM satellite schema and on the ESO release code, as the latest VVV and VMC submissions had been rejected for a variety of minor reasons. Non-survey Data Release: No new releases. Astrogrid deployment, VO & Data Analysis services: MSH noted that the VO services would need setting up for the recent VSA public releases. Miscellaneous: Nothing for the minutes. -- Scanned by iCritical.